# Treewidth of generalized Hamming graph, bipartite Kneser graph and generalized Petersen graph

### v1

@@ Abstract
Let $t,q$ and $n$ be positive integers. Write $[q]=\{1,2,\ldots,q\}$. The generalized Hamming graph $H(t,q,n)$ is the graph whose vertex set is the cartesian product of $n$ copies of $[q]$ $(q\geq 2)$, where two vertices are adjacent if their Hamming distance is at most $t$. In particular, $H(1,q,n)$ is the well-known Hamming graph and $H(1,2,n)$ is the hypercube. In 2006, Chandran and Kavitha described the asymptotic value of $tw(H(1,q,n))$, where $tw(G)$ denotes the treewidth of $G$. In this paper, we give the exact pathwidth of $H(t,2,n)$ and show that $tw(H(t,q,n))=\Theta(tq^{n}/\sqrt{n})$ when $n$ goes to infinity. Based on those results, we show that the treewidth of bipartite Kneser graph $BK(n,k)$ is $\binom{n}{k}-1$ when $n$ is sufficient large relative to $k$ and the bounds of $tw(BK(2k+1,k))$ are given. Moreover, we present the bounds of the treewidth of generalized Petersen graph.
@@ 1 Introduction
Treewidth is a well-studied parameter in graph theory. Many NP-complete problem can be solved in polynomial time on graphs of bounded treewidth [31, 6, 10]. Besides, treewidth is also useful in structural graph theory. For example, Robertson and Seymour used it to prove the Graph Minor Theorem [25, 26, 27]. In the past decades, there are lots of literatures investigating the treewidth of certain graphs (see for example, [14, 15, 19, 22, 30]). However, it is difficult to estimate the treewidth even asymptotically.
Throughout this paper, graphs are finite, simple and undirected. Let $V(G)$ and $E(G)$ denote the vertex set and edge set of $G$, respectively. The degree of a vertex $v\in V(G)$ in $G$ is denoted by $d_{G}(v)$, and we ignore the subscript $G$ in case of no ambiguity. The maximum and minimum degree of $G$ are denotes as $\Delta(G)$ and $\delta(G)$, respectively. Write $[n]=\{1,2,\ldots,n\}$ and denote by $\binom{[n]}{k}$ the family of all $k$-subsets of $[n]$. The treewidth and pathwidth of graph $G$ are denoted by $tw(G)$ and $pw(G)$, respectively.
The Hamming graph $H(q,n)$ is the graph on $q^{n}$ vertices, which correspond to all $n$-vectors whose components are from $[q]$ ($q\geq 2$). When $q=2$, it is more commonly to treat each vertex of $H(2,n)$ as a binary $n$-vector, that is, a $n$-vector whose components are from $\{0,1\}$. Two of the vertices in $H(q,n)$ are adjacent only when they differ in just one coordinate. The Hypercube graph is a special case of Hamming graph when $q=2$ which is well-studied in parallel computing, coding theory and many other areas[1, 16, 17]. The Hamming distance of two $n$-vectors is the number of coordinates that one differs from the other. Two vertices in a Hamming graph are adjacent if and only if their Hamming distance is no more than $1$. If two vertices in $H(q,n)$ is adjacent if and only if their Hamming distance is no more than $t$, then we denote the generalized Hamming graph as $H(t,q,n)$. Then $H(1,q,n)=H(q,n)$ and $H(1,2,n)$ is the hypercube graph. We call $H(t,2,n)$ the generalized hypercube graph and we will give the exact value of its pathwidth and bandwidth as Theorem 1.1.
@@
$pw(H(t,2,n))=bw(H(t,2,n))=\sum_{k=\left\lfloor(n-t)/2\right\rfloor}^{\left\lfloor(n-t)/2\right\rfloor+t-1}\binom{n}{k}+\sum_{a=0}^{\left\lfloor(n-t-1)/2\right\rfloor}\left(\binom{t+2a}{t+a-1}-\binom{t+2a}{a-1}\right).$ (1)
The treewidth of $H(q,n)$ is asymptotically $\Theta(q^{n}/\sqrt{n})$ from [8]. In this paper, we generalize this result to generalized Hamming graph $H(t,q,n)$ as Theorem 1.2. The previous result about $tw(H(q,n))$ can be derived from our result by letting $t=1$.
@@
Let $t$ and $q$ be positive integers. Then $tw(H(t,q,n))=\Theta(tq^{n}/\sqrt{n})$, when $n$ goes to infinity.
Kneser graph $K(n,k)$ is the graph whose vertex set is $\binom{[n]}{k}$ and where two vertices are adjacent if and only if the two corresponding sets are disjoint. The treewidth of Kneser graph is studied by Harvey in 2014 [14]. Generalized Kneser graph and $q$-Kneser graph are two derived class from Kneser graph and their treewidth can also be exactly described when $n$ is large enough [7, 23]. Bipartite Kneser graph is also an important variant of Kneser graph whose vertex set is $\binom{[n]}{k}$ and $\binom{[n]}{n-k}$ denoted by $BK(n,k)$ where $2k\leq n$. A $k$-subset $A$ and a $(n-k)$-subset $B$ in bipartite Kneser graph is adjacent if $A\subseteq B$. We will give the exact value of $tw(BK(n,k))$ when $n$ is sufficient large relative to $k$ as Theorem 1.3. When $n=2k+1$, we give the bounds of the treewidth of bipartite Kneser graph as Theorem 1.4.
@@
As $k$ goes to infinity,
$\Omega\left(\frac{1}{k}\binom{2k+1}{k}\right)=tw(BK(2k+1,k))\leq tw(J(2k+1,k))=O\left(\binom{2k+1}{k}\right).$
@@
$V(G_{n,k})$ =$=\{v_{1},\ldots,v_{n},u_{1},\ldots,u_{n}\},$
$E(G_{n,k})$ =$={\{v_{i}u_{i}\}}\cup{\{v_{i}v_{i+1}\}}\cup{\{u_{i}u_{i+k}\}},i=1,2,\ldots,n,$
@@
Let $n$ and $k$ be positive integers satisfying that $n\geq 8{(2k+2)}^{2}$. Then $22k+1\leq tw(G_{n,k})\leq pw(G_{n,k})\leq 2k+$ 2.
The rest of this paper is organized as follows. In Section 3, we give a proof of the treewidth of generalized Hamming graph. In Section 4, we study the treewidth of bipartite Kneser graph and Johnson graph. In Section 5, we study the treewidth of generalized Petersen graph.
@@ 2 Preliminaries and definitions
Let $X\subseteq V(G)$ be a subset of vertices and $G[X]$ be the subgraph induced in $G$ by $X$. Define $G-X=G[V(G)-X]$. Given $p\in(0,1)$, define the $p$-separator of $G$ to be a subset $X\subseteq V(G)$ such that no component of $G-X$ contains more than $p|V(G)-X|$ vertices. Proposition 2.4 describes the relationship between treewidth and separator.
@@
For any graph $G$, there exists a separator $X$ of $G$ with at most $tw(G)+1$ vertices. And there exists a partition of $V(G)-X$ into sets $A$ and $B$ such that
@@ 3.1 Bandwidth of Hamming graph $H(t,2,n)$
In this section, we intend to prove Theorem 1.1. When $t=1$, Eq 1 is exactly the same as Eq 3.
To prove Theorem 1.1, we need some preparation. In [13], Harper showed that if an ordering $\varphi$ of $G(V,E)$ is in Hales order (i.e., a Hales numbering), then $\max\limits_{e\in E}\Delta(e,\varphi)$ takes minimum over all numbering, that is, $bw(G)=\min\limits_{\phi}\max\limits_{e\in E}\Delta(e,\phi)=\max\limits_{e\in E}\Delta(e,\varphi)$. Harper [13] also give a sample of Hales numbering of $H(t,2,n)$ [^1] It is not explicitly stated in [13]. Harper first gave a Hales numbering of hypercube and, then claimed that the numbering is also in Hales order for the distance generalized graph (that is, $H(t,2,n)$) in comments.. From [29], we can build such Hales numbering $\varphi$ in the following way.
@@
where $\mathbf{1}$ (resp. $\mathbf{0}$) is an all one (resp. all zero) $\binom{n-1}{k-1}$ (resp. $\binom{n-1}{k}$) column vector. Clearly, the rows of $A_{k}^{(n)}$ enumerates all binary vectors of length $n$ whose number of ‘$1$’ is $k$. Define $S^{(n)}$ as following:
@@
Then $S^{(n)}$ is an $2^{n}\times n$ matrix. Each row of $S^{(n)}$ excatly corresponds to a $n$-vector (a vertices of $H(t,2,n)$) and vice versa. Let $\eta^{(n)}$ be the ordering of $H(t,2,n)$ defined by row order of $S^{(n)}$, that is, if $v\in V(H(t,2,n))$ corresponds to the $i$-th row of $S^{(n)}$, then let $\eta^{(n)}(v)=i$.
$\eta^{(n)}$ is a Hales numbering of $H(t,2,n)$ (see [29] and [13]) and, hence, we have the following Proposition 3.2.
@@
[…]
@@
Considering $M^{(t,n)}_{k_{1},k_{2}}$ with $k_{1}\leq k_{2}\leq k_{1}+t$, it is easy to verify that if the parity of $k_{2}-k_{1}$ and $t$ is different, then $M^{(t,n)}_{k_{1},k_{2}}=M^{(t-1,n)}_{k_{1},k_{2}}$. Hence, we only need to calculate all $M^{(t,n)}_{k_{1},k_{2}}$ with $k_{1}\leq k_{2}\leq k_{1}+t$ and $k_{2}-k_{1}\equiv t\pmod{2}$. In this case, assume $t-(k_{2}-k_{1})=2s$. Then $M^{(t,n)}_{k,k+t-2s}$ is non-zero. If $t\geq n-1$, then $M^{(t,n)}_{k,k+t-2s}$ is an all-one matrix (except the main diagonal when $t=2s$) and, thus, $r(M^{(t,n)}_{k,k+t-2s})=\binom{n}{k}+\binom{n}{k+t-2s}-1$. When $11\leq t<n-$ 1, we let
@@
$C_{t,n,k,s}=\sum\limits_{m=n-s+1}^{n}\binom{m-1}{k+t-2s-1},$
@@
Then we have the following results.
@@
Let $t,n,k,s$ ben non-negative integers satisfy that $t\geq 2s,n\geq 1,t\geq 1$ and $k+t-2s\leq n$. Then
$r(M^{(t,n)}_{k,k+t-2s})$ =$=\left\{\begin{array}[]{ll}\binom{n}{k}+\binom{n}{k+t-2s}-1&\mbox{if }t\geq n-1,\\ \binom{n}{k}+A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}+C_{t,n,k,s}&\mbox{if }t<n-1,0\leq k-s\leq\left\lfloor(n-t)/2\right\rfloor,\\ \binom{n}{k}+B_{t,n,k,s}+C_{t,n,k,s}&\mbox{if }t<n-1,\left\lfloor(n-t)/2\right\rfloor\leq k-s\leq n-t,\\ \binom{n}{k}+D_{t,n,k,s}&\mbox{if $t<n-1$, $k-s\leq 0$ or $k-s\geq n-t$.}\end{array}\right.$ (13)
@@
From Eq 10 and 12, we can prove Lemmas 3.10 and 3.11 by induction. The complete proofs of Lemmas 3.10 and 3.11 can be find in Appendix A and Appendix B, respectively. Lemmas 3.11 and 3.6 show the exact bandwidth of $H(t,2,n)$.
@@
We have $pw(H(t,2,n))$ $=bw(H(t,2,n))$.
Proof.  By Theorem 3.12, we have $pw(H(t,2,n))\geq b_{v}(s,H(t,2,n))$ for all $1\leq s\leq 2^{n}$. Then $pw(H(t,2,n))\geq\max\limits_{1\leq s\leq 2^{n}}b_{v}(s,H(t,2,n))=bw(H(t,2,n))$. Combining Eq 2, we have $pw(H(t,2,n))$ $=bw(H(t,2,n))$. $\square$
Theorem 1.1 can be derived from Lemmas 3.63.11 and 3.13.

### v3

@@ Abstract
Let $t,q$ and $n$ be positive integers. Write $q]=\{1,2,\ldots,q\}$. The generalized Hamming graph $H(t,q,n)$ is the graph whose vertex set is the cartesian product of $n$ copies of $[q]$ $(q\geq 2)$, where two vertices are adjacent if their Hamming distance is at most $t$. In particular, $H(1,q,n)$ is the well-known Hamming graph and $H(1,2,n)$ is the hypercube. In 2006, Chandran and Kavitha described the asymptotic value of $tw(H(1,q,n))$, where $tw(G)$ denotes the treewidth of $G$. In this paper, we give the exact pathwidth of $H(t,2,n)$ and show that $tw(H(t,q,n))=\Theta(tq^{n}/\sqrt{n})$ when $n$ goes to infinity. Based on those results, we show that the treewidth of the bipartite Kneser graph $BK(n,k)$ is $\binom{n}{k}-1$ when $n$ is sufficient large relative to $k$ and the bounds of $tw(BK(2k+1,k))$ are given. Moreover, we present the bounds of the treewidth of generalized Petersen graph.
@@ 1 Introduction
Treewidth is a well-studied parameter in graph theory. Many NP-complete problem can be solved in polynomial time on graphs of bounded treewidth [36, 7, 11]. Besides, treewidth is also useful in structural graph theory. For example, Robertson and Seymour used it to prove the Graph Minor Theorem [30, 31, 32]. In the past few decades, there has been much literature investigating the treewidth of certain graphs (see for example, [17, 18, 22, 25, 35]). However, it is difficult to estimate the treewidth even asymptotically.
Throughout this paper, graphs are finite, simple and undirected. Let $V(G)$ and $E(G)$ denote the vertex set and edge set of $G$, respectively. The degree of a vertex $v\in V(G)$ in $G$ is denoted by $d_{G}(v)$, and we ignore the subscript $G$ in case of no ambiguity. The maximum and minimum degree of $G$ are denoted as $\Delta(G)$ and $\delta(G)$, respectively. Write $n]=\{1,2,\ldots,n\}$ and denote by $\binom{[n]}{k}$ the family of all $k$-subsets of $[n]$. The treewidth, pathwidth and bandwidth of graph $G$ are denoted by $tw(G)$, $pw(G)$ and $bw(G)$, respectively. It is worth mentioning that the isoperimetric problem of the hypercube is an important topic which provides a foundation for our subsequent proof. The readers may refer to [15, 14, 29, 16, 3] for more details.
The Hamming graph $H(q,n)$ is the graph on $q^{n}$ vertices, which correspond to all $n$-vectors whose components are from $[q]$ ($q\geq 2$). When $q=2$, it is more common to treat each vertex of $H(2,n)$ as a binary $n$-vector, that is, an $n$-vector whose components are from $\{0,1\}$. Two of the vertices in $H(q,n)$ are adjacent only when they differ in just one coordinate. The hypercube graph is a special case of the Hamming graph when $q=2$ which is well-studied in parallel computing, coding theory and many other areas[1, 19, 20].
The Hamming distance of two $n$-vectors is the number of coordinates where one differs from the other. Two vertices in a Hamming graph are adjacent if and only if their Hamming distance is no more than $1$. If two vertices in $H(q,n)$ is adjacent if and only if their Hamming distance is no more than $t$, then we denote the generalized Hamming graph as $H(t,q,n)$. Then $H(1,q,n)=H(q,n)$ and $H(1,2,n)$ is the hypercube graph. We call $H(t,2,n)$ the generalized hypercube graph and we will give the exact value of its pathwidth and bandwidth as Theorem 1.1. When $t=1$, our result covers the previous results of hypercube graph. It is interesting to see that in the generalized form of hypercube graph, its treewidth still can be described by an exact expression. Having exact form of treewidth is not a common phenomenon, especially with such a complicated form.
@@
$pw(H(t,2,n))=bw(H(t,2,n))$ (1)
$=$ $\sum_{k=\left\lfloor(n-t)/2\right\rfloor}^{\left\lfloor(n-t)/2\right\rfloor+t-1}\binom{n}{k}+\sum_{a=0}^{\left\lfloor(n-t-1)/2\right\rfloor}\left(\binom{t+2a}{t+a-1}-\binom{t+2a}{a-1}\right)$.
The treewidth of $H(q,n)$ is asymptotically $\Theta(q^{n}/\sqrt{n})$ from [9]. To be more specific, there exists constants $c_{1}$ and $c_{2}$ not depending on $q$, such that for sufficiently large $n$, $c_{1}q^{n}/\sqrt{n}\leq tw(H(q,n))\leq c_{2}q^{n}/\sqrt{n}$. In this paper, we generalize this result to generalized Hamming graph $H(t,q,n)$ as Theorem 1.2. The previous result about $tw(H(q,n))$ can be derived from our result by letting $t=1$. It is interesting to see that the distance variable $t$ in the generalized Hamming graph $H(t,q,n)$ has linear impact on its treewidth.
@@
There exists constant $c_{1}$ and $c_{2}$ not depending on $t$ or $q$, such that for any positive integers $t$ and $q$. When $n$ is sufficiently large, we have that
$c_{1}tq^{n}/\sqrt{n}\leq tw(H(t,q,n))\leq c_{2}tq^{n}/\sqrt{n}.$
Kneser graph $K(n,k)$ is the graph whose vertex set is $\binom{[n]}{k}$ and where two vertices are adjacent if and only if the two corresponding sets are disjoint. The treewidth of Kneser graph is studied by Harvey in 2014 [17]. Generalized Kneser graph, $q$-Kneser graph and generalized $q$-Kneser graph are three derived class from Kneser graph and their treewidth can also be exactly described when $n$ is large enough [8, 26, 27]. Bipartite Kneser graph is also an important variant of Kneser graph whose vertex set is $\binom{[n]}{k}$ and $\binom{[n]}{n-k}$ denoted by $BK(n,k)$ where $2k\leq n$. A $k$-subset AA and a (nk)(n-k)-subset $B$ in bipartite Kneser graph is adjacent if $A\subseteq B$. We will give the exact value of $tw(BK(n,k))$ when $n$ is sufficient large relative to $k$ as Theorem 1.3. When $n=2k+1$, we give the bounds of the treewidth of the bipartite Kneser graph as Theorem 1.4.
@@
There exists two constants $c_{1}$ and $c_{2}$ such that for any positive integer $k$,
$c_{1}\frac{1}{k}\binom{2k+1}{k}\leq tw(BK(2k+1,k))\leq tw(J(2k+1,k))\leq c_{2}\binom{2k+1}{k}$.
@@
$V(G_{n,k})$ $=\{v_{1},\ldots,v_{n},u_{1},\ldots,u_{n}\},$
$E(G_{n,k})$ $={\{v_{i}u_{i}\}}\cup{\{v_{i}v_{i+1}\}}\cup{\{u_{i}u_{i+k}\}},i=1,2,\ldots,n,$
@@
Let $n$ and $k$ be positive integers satisfying that $n\geq 8{(2k+2)}^{2}$. Then we have
$2k+1\leq tw(G_{n,k})\leq pw(G_{n,k})\leq 2k+2.$
Theorem 1.31.4 and 1.5 can be viewed as different generalizations of Kneser graph. And in Theorem 1.4, it is interesting to see that the treewidth of the bipartite Kneser graph has a close relationship with that of the well-studied Johnson graph. The treewidth of Johnson graph $J(n,k)$ is also an interesting topic. When $k=2$, $tw(J(n,2))$ and $pw(J(n,2))$ have exact formulas while for other $k$ it remains unknown [12]. Our result may help to have a deeper understanding of the treewidth of Johnson graph.
The rest of this paper is organized as follows. In Section 3, we give a proof of the treewidth of generalized Hamming graph (Theorem 1.1 and Theorem 1.2). In this part, we mainly use properties of bandwidth and techniques of Hales numbering and the isoperimetric problems by Harper [15, 14].
In Section 4, we study the treewidth of the bipartite Kneser graph and Johnson graph (Theorem 1.3 and Theorem 1.4). In this part, we mainly use the techniques of separators, properties of cross-intersecting families and chordal completions. Since the Johnson graph can be viewed as a slice of generalized Hamming graph, we need the results in Section 3 to prove our results.
In Section 5, we study the treewidth of generalized Petersen graph (Theorem 1.5). In this part, we mainly use brambles and path-decomposition constructions.
@@ 2 Preliminaries and definitions
Let $X\subseteq V(G)$ be a subset of vertices and $G[X]$ be the subgraph induced in $G$ by $X$. Define GX=G[V(G)X]G-X=G[V(G)-X]. Given $p\in(0,1)$, define the $p$-separator of $G$ to be a subset $X\subseteq V(G)$ such that no component of GXG-X contains more than p|V(G)X|p|V(G)-X| vertices. Proposition 2.4 describes the relationship between treewidth and separator.
@@
For any graph $G$, there exists a separator $X$ of $G$ with at most $tw(G)+1$ vertices. And there exists a partition of V(G)XV(G)-X into sets $A$ and $B$ such that
@@ 3.1 Bandwidth of Hamming graph $H(t,2,n)$
In this section, we intend to prove Theorem 1.1. When $t=1$, Eq 1 is exactly the same as Eq 3. Thus, we can view Theorem 1.1 as a generalization of Proposition 3.1. Using the following techniques, we will derive some recursion formulas (Proposition 3.9 and Proposition 3.10) and use induction to prove Theorem 1.1. However, Eq 1 is much more complicated and once we know its formula, it is always “easy” to verify it when we have some recursion formulas using induction. The exact expression of Eq 10 and 12 actually come from our elegant observation from some instances with computer assistance.
To prove Theorem 1.1, we need some preparation. In [15], Harper showed that if an ordering $\varphi$ of $G(V,E)$ is in Hales order (i.e., a Hales numbering), then $\max\limits_{e\in E}\Delta(e,\varphi)$ takes minimum over all numbering, that is, $bw(G)=\min\limits_{\phi}\max\limits_{e\in E}\Delta(e,\phi)=\max\limits_{e\in E}\Delta(e,\varphi)$.
Definition 3.2 (Hales numbering [15])
If there exists an ordering $\phi$ all of whose beginning segments obey the following two conditions, then we call such orderings Hales numberings. Note that Hales numbering does not always exists and it is not unique.
  1. 1.
    For a set of $l$ vertices, let $\Phi_{l}$ be the number of vertices in the set having neighbors not in the set. $\Phi_{l}$ must be minimized for all beginning segments $S_{l}=\{v\in V\mid\phi(v)\leq l\}$.
  2. 2.
    The $\Phi_{l}^{\prime}=l-\Phi_{l}$ “interior vertices” of $S_{l}$ must be numbered $1,2,\ldots,\Phi_{l}^{\prime}$, i.e., have the lowest possible numbers on them.
Harper [15] also give a sample of Hales numbering of H(t,2,n)H(t,2,n) [^1] In [15], Harper did not prove this statement. Harper first gave a Hales numbering of hypercube and, then claimed that the numbering is also in Hales order for the distance generalized graph (that is, $H(t,2,n)$) in comments (III (b)).. From [34], we can build such Hales numbering $\varphi$ in the following way.
@@
[…]
@@
Then $S^{(n)}$ is an $2^{n}\times n$ matrix. Each row of $S^{(n)}$ excatly corresponds to an $n$-vector (a vertices of $H(t,2,n)$) and vice versa. Let $\eta^{(n)}$ be the ordering of $H(t,2,n)$ defined by row order of $S^{(n)}$, that is, if $v\in V(H(t,2,n))$ corresponds to the $i$-th row of $S^{(n)}$, then let $\eta^{(n)}(v)=i$.
$\eta^{(n)}$ is a Hales numbering of $H(t,2,n)$ (see [34] and [15]) and, hence, we have the following Proposition $3.$
@@
$M(\eta_{1}(u),\eta_{2}(v))=\left\{\begin{array}[]{cc}1&\mbox{if $\{u,v\}\in E$},\\ 0&\mbox{otherwise}.\end{array}\right$.
@@
Considering $M^{(t,n)}_{k_{1},k_{2}}$ with $k_{1}\leq k_{2}\leq k_{1}+t$, it is easy to verify that if the parity of $k_{2}-k_{1}$ and $t$ is different, then $M^{(t,n)}_{k_{1},k_{2}}=M^{(t-1,n)}_{k_{1},k_{2}}$. Hence, we only need to calculate all $M^{(t,n)}_{k_{1},k_{2}}$ with $k_{1}\leq k_{2}\leq k_{1}+t$ and $k_{2}-k_{1}\equiv t\pmod{2}$. In this case, assume $t-(k_{2}-k_{1})=2s$. Then $M^{(t,n)}_{k,k+t-2s}$ is non-zero. If $t\geq n-1$, then $M^{(t,n)}_{k,k+t-2s}$ is an all-one matrix (except the main diagonal when $t=2s$) and, thus, $r(M^{(t,n)}_{k,k+t-2s})=\binom{n}{k}+\binom{n}{k+t-2s}-1$. When $1\leq t<n-1$, we let
@@
$C_{t,n,k,s}=\sum\limits_{m=n-s+1}^{n}\binom{m-1}{k+t-2s-1}=\binom{n}{k+t-2s}-\binom{n-s}{k+t-2s}$,
@@
Here we define $\binom{x}{y}=0$ if $y<0$ or $y>x$. Note that all $A^{(1)},A^{(2)},B,C,D$ terms are non-negative. It is easy to verify that if k+t2s=nk+t-2s=n, then $D_{t,n,k,s}=0$, otherwise k+t2s<nk+t-2s<n, then $D_{t,n,k,s}=\binom{n}{k+t-2s}-\binom{k+t-2s}{k+t-2s}=\binom{n}{k+t-2s}-1$. Then we have the following results.
@@
Let $t,n,k,s$ be non-negative integers satisfy that $t\geq 2s,n\geq 1,t\geq 1$ and $k+t-2s\leq n$. Then
$r(M^{(t,n)}_{k,k+t-2s})$ $=\left\{\begin{array}[]{ll}\binom{n}{k}+\binom{n}{k+t-2s}-1&\mbox{if }t\geq n-1,\\ \binom{n}{k}+A^{(1)}_{t,n,k,s}+A^{(2)}_{t,n,k,s}+C_{t,n,k,s}&\mbox{if }t<n-1,0\leq k-s\leq\left\lfloor(n-t)/2\right\rfloor,\\ \binom{n}{k}+B_{t,n,k,s}+C_{t,n,k,s}&\mbox{if }t<n-1,\left\lfloor(n-t)/2\right\rfloor\leq k-s\leq n-t,\\ \binom{n}{k}+D_{t,n,k,s}&\mbox{if $t<n-1$, $k-s\leq 0$ or $k-s\geq n-t$.}\end{array}\right$. (13)
@@
From Eq 10 and 12, we can prove Lemmas 3.11 and 3.12 by induction. The complete proofs of Lemmas 3.11 and 3.12 can be find in Appendix A and Appendix B, respectively. Lemmas 3.12 and 3.7 show the exact bandwidth of $H(t,2,n)$.
@@
We have $pw(H(t,2,n))$ =bw(H(t,2,n))=bw(H(t,2,n)).
Proof.    By Theorem 3.13, we have $pw(H(t,2,n))\geq b_{v}(s,H(t,2,n))$ for all $1\leq s\leq 2^{n}$. Then $pw(H(t,2,n))\geq\max\limits_{1\leq s\leq 2^{n}}b_{v}(s,H(t,2,n))=bw(H(t,2,n))$. Combining Eq 2, we have $pw(H(t,2,n))$ =bw(H(t,2,n))=bw(H(t,2,n)). $\square$
Theorem 1.1 can be derived from Lemmas 3.73.12 and 3.14.